# Evaluating the evaluators: Towards human-aligned metrics for missing markers reconstruction

### v1

@@ Abstract.
Animation data is often obtained through optical motion capture systems, which utilize a multitude of cameras to establish the position of optical markers. However, system errors or occlusions can result in missing markers, the manual cleaning of which can be time-consuming. This has sparked interest in machine learning-based solutions for missing marker reconstruction in the academic community. Most academic papers utilize a simplistic mean square error as the main metric. In this paper, we show that this metric does not correlate with subjective perception of the fill quality. We introduce and evaluate a set of better-correlated metrics that can drive progress in the field.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Motion capture (MoCap) technology is extensively utilized in the film and video game industry to animate virtual characters with lifelike movements. Even though marker-based solutions capture body movements with more precision than purely video-based trackers, even state-of-the-art MoCap equipment and software suffer from occasional marker occlusion and mislabeling, as well as positional errors. Manual clean-up is therefore still necessary. As this process can be particularly time-consuming, even for experienced animators, the idea of automating MoCap data cleaning has long been of interest (Liu and McMillan, 2006). Since MoCap data cleaning can be formulated as a mapping problem from raw to clean data, machine learning (ML) solutions, and in particular deep learning methods, have been proposed to solve this problem, e.g. (Holden, 2018; Perepichka et al., 2019; Kaufmann et al., 2020; Cui and Sun, 2021; Xia et al., 2022; Lee et al., 2024).
@@
First, papers vary in terms of datasets and evaluation protocol. The general setup of related papers is to define an artificial distribution over missing markers and positional errors and to corrupt both the training and the testing data using this distribution (Holden, 2018). Approaches vary in terms of distributions of missing markers during test time. While some assume individual markers to be missing independently from each other (Perepichka et al., 2019; Chen et al., 2021; Yang et al., 2024), others assume gaps over different time windows (Kucherenko et al., 2018; Li et al., 2021; Skurowski and Pawlyta, 2024) or body parts (Cui and Sun, 2021; Lohit et al., 2021; Perepichka et al., 2019). The performance measured in these papers thus only indicates how well the model would perform on data corruptions that are distributed similar to the artificial distribution, not how well it would perform on real-world MoCap recordings (Pan et al., 2023). While we train our models on artificially corrupted data, we test on real-world raw data. As a result, we are able to assess model performance in a realistic setting.
@@
The second problem of current evaluation practices is that metrics focus on reconstruction errors and not on aspects that are important to the end user of the data, in this case animators. The predominant metric in related work is the mean squared error (MSE) between the systems predictions for missing markers and the ground truth position (Holden, 2018; Perepichka et al., 2019; Xia et al., 2022; Cui et al., 2019; Lohit et al., 2021; Li et al., 2021; Kieu et al., 2022; Cai et al., 2021; Yang et al., 2024). However, MSE-like metrics fail to measure spatial and temporal smoothness. Both temporal consistency and consistent bodily proportions are important to animators because they make the solving to a skeleton and manual editing easier. For this reason, spatial and temporal consistency are more important than the distance between the predicted sequence and a ground truth sequence. We, therefore, propose metrics that align with animators’ needs and assess how well different metrics correlate with domain expert judgment of reconstruction quality.
@@
In more detail, our hypothesis is that metrics should take spatial and temporal consistency into account. To verify this hypothesis, we propose and evaluate such metrics, namely bone distance preservation (BDP) and velocity distance (VD). To see if these metrics align with the perceptual quality of quality, we correlate metric values with human judgment of animation quality. To this end, we perform a user study in which domain experts are asked to rate motion capture sequences after they have been cleaned by different ML systems and solved onto a skeleton. We then perform a correlation analysis to understand the correlation between human judgment and different metrics.
It is important to note that this work focuses on the evaluation of ML-based MoCap cleaning systems and not on presenting a novel, state-of-the-art model. We do, therefore, not benchmark our models on publicly available datasets, which do not provide matching raw and cleaned data. Instead, we use an in-house dataset of paired raw and clean data so that we can establish the validity of metrics in a real-world setting. Our dataset poses a particular challenge as it consists of two actors performing different grappling motions. Due to occlusions by the other actor, it does, therefore, contain large gaps. On average, gaps consist of 25 % of markers for on average of 1 second.
In conclusion, our contributions are as follows:
  • Inspired by domain expert feedback, we propose a set of new metrics to measure the performance of missing marker solutions.
  • To evaluate the validity of different metrics, we design and carry out a human subject study that measures perceived quality of the predicted markers.
@@ 2. Related work
This paper primarily focuses on evaluation practices. Therefore we only briefly touch upon different modeling approaches to solve the missing marker problem (Section 2.1) and dedicate more attention to metrics (Section 2.2). For a detailed review of the missing markers reconstruction field we refer the reader to Martini et al. (2023).
@@ 3.1. Problem formulation
Let $Y\in\mathcal{R}^{(T,N_{m})}$ be a matrix that represents a sequence of $T$ frames and a skeleton that consists of $N_{m}$ markers. We distinguish between raw observations $Y_{raw}$ and cleaned observations $Y_{clean}$. Further, we use dot $\dot{Y}$ to denote a sequence that has been transformed by noise and hat $\hat{Y}$ to denote a predicted sequence.
The usual problem statement of missing markers reconstruction problem is to assume a ground truth marker sequence $Y_{clean}$. In order to simulate marker occlusion and positional noise, $Y$ is transformed with the help of a masking matrix $O\in[0,1]^{(T,N_{m})}$ and a noise matrix $\Sigma\in\mathcal{R}^{(T,N_{m})}$ as follows
@@
We then train the model to denoise the marker values that were randomly shifted. During testing, instead of filling the missing marker sequences with zeros, we use linear interpolation between the last observation of a marker and the first observation following the gap:
@@ 4.1. Dataset
Our internal dataset contains recordings of two actors emulating various close-combat actions, mainly grappling. Each actor has 63 markers, for a total of 126 markers. The data was recorded at 120 frames per second (fps). In total, we had access to 230,365 frames meaning roughly 30 minutes of motion data. Those were split in the following way: 217,318 for training, 6,263 for validation, and 6,784 for testing.
@@ 4.2.2. Bone distance preservation
BDP with GT is defined as:
@@
BDP without GT is then defined as:
@@ 4.2.3. Velocity distance
To measure temporal coherence we propose to calculate a velocity distance (VD). As in the case of bone distance preservation, there are two ways to define a velocity distance: either with respect to the ground truth or with respect to the previous time frame. We will call the two VD with GT and VD without GT.
@@
(11) $\sqrt{\frac{1}{T}\sum_{t=2}^{N}\lVert v_{t}-\hat{v}_{t}\rVert^{2}}.$
@@ Acknowledgements.
We are grateful to Tomas Tjarnberg and Alvaleia Thorsson for their data collection efforts on this project; to Harald Zlattinger and Chris Battson for their invaluable support, guidance, and resources; to Timur Solovev for his engineering support; to Jim Royal for help with crafting visuals; and to Clement Lescalet, Jeremy Mathiesen, Aaron Punga, John Aguilar, Jae Lee, Kristy Clarke, Sam Kwok, Audrey Ito, Skyla Li, Vince Ng, Marcelo Sakai, Brennan Robinson, Vincent Hung, and Lucy Guo for participating in the user study.

### v2

@@ Abstract.
Animation data is often obtained through optical motion capture systems, which utilize a multitude of cameras to establish the position of optical markers. However, system errors or occlusions can result in missing markers, the manual cleaning of which can be time-consuming. This has sparked interest in machine learning-based solutions for missing marker reconstruction in the academic community. Most academic papers utilize a simplistic mean square error as the main metric. In this paper, we show that this metric does not correlate with subjective perception of the fill quality. Additionally, we introduce and evaluate a set of better-correlated metrics that can drive progress in the field.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Conference; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Optical motion capture (MoCap) is a process that utilizes a multitude of cameras to establish the position of optical markers. This process is followed by fitting a skeleton into the point cloud of optical markers’ positions. The latter operation is commonly referred to as solving. MoCap technology is extensively utilized in the film and video game industry to animate virtual characters with lifelike movements. Although marker-based solutions capture body movements with more precision than purely video-based trackers, even state-of-the-art MoCap equipment and software suffer from occasional marker occlusion and mislabeling, as well as positional errors. Manual clean-up is therefore still necessary. As this process can be particularly time-consuming, even for experienced animators, the idea of automating MoCap data cleaning has long been of interest (Liu and McMillan, 2006). Machine learning (ML) solutions, and in particular deep learning methods, have been proposed to solve this problem, e.g. (Holden, 2018; Perepichka et al., 2019; Kaufmann et al., 2020; Cui and Sun, 2021; Xia et al., 2022; Lee et al., 2024).
@@
First, papers vary in terms of datasets and evaluation protocol. The general setup of related papers is to define an artificial distribution over missing markers and positional errors and to corrupt both the training and the testing data using this distribution (Holden, 2018). Approaches vary in terms of distributions of missing markers during test time. While some assume individual markers to be missing independently from each other (Perepichka et al., 2019; Chen et al., 2021; Yang et al., 2024), others assume gaps over different time windows (Kucherenko et al., 2018; Li et al., 2021; Skurowski and Pawlyta, 2024) or body parts (Cui and Sun, 2021; Lohit et al., 2021; Perepichka et al., 2019). The performance measured in these papers thus only indicates how well the model would perform on data corruptions that are distributed similar to the artificial distribution, not how well it would perform on real-world MoCap recordings (Pan et al., 2023). We evaluate models using real-world data to assess their performance in a realistic setting.
@@
The second problem of current evaluation practices is that metrics focus on reconstruction errors and not on aspects that are important to the end user of the data, in this case animators. The predominant metric in related work is the mean squared error (MSE) between the systems predictions for missing markers and the ground truth position (Holden, 2018; Perepichka et al., 2019; Xia et al., 2022; Cui et al., 2019; Lohit et al., 2021; Li et al., 2021; Kieu et al., 2022; Cai et al., 2021; Yang et al., 2024). However, MSE-like metrics fail to measure spatial and temporal correlation. Both temporal consistency and consistent bodily proportions are important to animators because they make the solving to a skeleton and manual editing easier. For this reason, spatial and temporal consistency are more important than the distance between the predicted sequence and a ground truth sequence. We, therefore, propose metrics that align with animators’ needs and assess how well different metrics correlate with domain expert judgment of reconstruction quality.
@@
To verify that metrics should take spatial and temporal consistency into account, we propose and evaluate such metrics, namely Bone Distance Preservation (BDP) and Velocity Distance (VD). To see if these metrics align with the perceptual quality, we correlate metric values with human judgment of animation quality. To this end, we perform a user study in which domain experts are asked to rate motion capture sequences after they have been cleaned by different ML systems and solved onto a skeleton. We then analyse correlations between human judgment and different metrics.
It is important to note that this work focuses on the evaluation of ML-based MoCap cleaning systems and not on presenting a novel, state-of-the-art system. Therefore, we do not benchmark our models on publicly available datasets.
Furthermore, while we do propose new metrics and modifications of existing metrics, the main point of this paper is not to provide a solution but rather to outline a problem with the commonly used metrics in the field.
In conclusion, our contributions are:
  • Conducting the first human-subject study to assess the perceived quality of predicted markers.
  • Developing new metrics to evaluate missing marker reconstruction solutions, including those that do not require access to ground truth data.
  • Performing a correlation analysis between subjective and objective ratings to validate various metrics.
@@ 2. Related work
This paper primarily focuses on evaluation practices. Therefore we only briefly touch upon different modeling approaches to solve the missing marker problem (Section 2.1) and dedicate more attention to metrics (Section 2.2). For a detailed review of the missing markers reconstruction field we refer the reader to (Martini et al., 2023).
@@ 2.2. Metrics
One way to take temporal cohesion into account is to calculate velocity distance. This has not been done before in the field of missing markers reconstruction, to the best of our knowledge, and we propose two novel metrics based on velocity distance which are detailed in Section 4.2.
To incorporate temporal cohesion, we propose using the velocity distance. We suggest two novel metrics based on velocity distance, detailed in Section 4.2.
@@ 3.1. Problem formulation
Let $Y\in\mathcal{R}^{(T,3N_{m})}$ be a matrix that represents a sequence of $T$ frames and a skeleton that consists of $N_{m}$ markers. We distinguish between raw observations $Y_{raw}$ and cleaned observations $Y_{clean}$. Further, we use dot $\dot{Y}$ to denote a sequence that has been transformed by noise and hat $\hat{Y}$ to denote a predicted sequence.
The usual problem statement of missing markers reconstruction problem is to assume a ground truth marker sequence $Y_{clean}$. In order to simulate marker occlusion and positional noise, $Y$ is transformed with the help of a masking matrix $O\in[0,1]^{(T,3N_{m})}$ and a noise matrix $\Sigma\in\mathcal{R}^{(T,3N_{m})}$ as follows
@@
We then train the model to denoise the marker values that were randomly shifted. During testing, instead of filling the missing marker sequences with zeros, we use cubic interpolation between the last observation of a marker and the first observation following the gap:
@@ 4.1. Dataset
Our internal dataset contains recordings of two actors emulating various close-combat actions, mainly grappling. Due to occlusions by the other actor, it does, therefore, contain large gaps. On average, gaps consist of 25 % of markers for on average of 1 second. Each actor has 63 markers, for a total of 126 markers. The data was recorded at 120 frames per second (fps). In total, we had access to 230,365 frames meaning roughly 30 minutes of motion data. Those were split in the following way: 217,318 for training, 6,263 for validation, and 6,784 for testing.
@@ 4.2.2. Bone distance preservation
BDP with GT was previously used (Cui et al., 2019; Kaufmann et al., 2020; Yin et al., 2021) and is defined as:
@@
BDP without GT was not used before and is defined as:
@@ 4.2.3. Velocity distance
To measure temporal coherence, we propose to calculate a velocity distance (VD). This has been previously used in other animation tasks (karras2017audio; holden2020learned), but not for missing marker reconstruction. As in the case of bone distance preservation, there are two ways to define a velocity distance: either with respect to the ground truth or with respect to the previous time frame. We will call the two VD with GT and VD without GT.
@@
(11) $\sqrt{\frac{1}{T}\sum_{t=2}^{T}\lVert v_{t}-\hat{v}_{t}\rVert^{2}}.$
@@ 8. Limitations
Finally, the dataset used in this work is proprietary, and we cannot publicly release it, which limits reproducibility.
@@ Acknowledgements.
We are grateful to Tomas Tjarnberg and Alvaleia Thorsson for their data collection efforts on this project; to Harald Zlattinger and Chris Battson for their invaluable support, guidance, and resources; to Timur Solovev for his engineering support; to Jim Royal for help with crafting visuals; to Mónica Villanueva Aylagas, Héctor Leon, and Alessandro Sestini for feedback on the paper; and to Clement Lescalet, Jeremy Mathiesen, Aaron Punga, John Aguilar, Jae Lee, Kristy Clarke, Sam Kwok, Audrey Ito, Skyla Li, Vince Ng, Marcelo Sakai, Brennan Robinson, Vincent Hung, and Lucy Guo for participating in the user study.